OpenStudy (anonymous):

Write an equation of a line that passes through the point (7, 3) and is parallel to the line y = -2/3x+3.

OpenStudy (anonymous):

a.y=3/2 x -3 b.3/2 x +3 c.-2/3 x +23/3 d. -2/3 x -23/3

OpenStudy (anonymous):

m = -2/3 y-3 = -2/3(x-7) y = -2/3x+14/3+3 y = -2/3+23/3 c.

OpenStudy (anonymous):

C again

OpenStudy (anonymous):

Well, yeah.

OpenStudy (anonymous):

ur the best

OpenStudy (anonymous):

i got it right

OpenStudy (anonymous):

Thanks.. :)))

OpenStudy (anonymous):

can you help me sommore

OpenStudy (anonymous):

The volume of a cone that fits exactly inside a cylinder is 3 cubic feet. What is the volume of the cylinder? (5 points) Select one: a. 6 cubic feet b. 9 cubic feet c. 12 cubic feet d. 18 cubic feet here